When the tension on a wire is 4N its length is l1. When the tension on the wire is 5N its length is l2. Find its natural length

Grade:11

2 Answers

Vikas TU
7 years ago
let x1 and x2 be the extension with 4N and 5N respectively.
then the eqns. would be:
4 = k * x1  and  5  = k * x2
and let l be the natural length of the wire.
Hence,
x1 + l = l1
and
x2 + l = l2 
put x1 and x2 values in Tensions after dividing them respectively:
Therefore,
4/5 = x1/x2 = > (l1 – l)/(l2 – l)
4l2 – 4l = 5l1 – 5l
l = 5l1 – 4l2
hence would be the nautral length.
Roshan agrawal
6 years ago
L = (T2l1 - T1l2)/T2 - T1 Where T2 and T1 = tension l1 and l2 = length This can be solved by equate the modules of elasticity that young`s modules